Proof using Osborne's rule that cosh2A=Cosh^2A+Sinh^2A

1
Expert's answer
2022-06-06T05:38:06-0400

The rule states that one replaces every occurrence of sine or cosine with the corresponding hyperbolic sine or cosine, and wherever one has a product of two sines, the product of the hyperbolic sines must be negated.


"\cos(2A)=\cos(A+A)=\cos A\cos A-\sin A\sin A"

By the Osborne's rule


"\cosh(2A)=\cosh(A+A)"

"=\cosh A\cosh A+\sinh A\sinh A"

Then


"\cosh(2A)=\cosh^2 A+\sinh^2 A"